# 3D-Guided Scalable Flow Matching for Generating Volumetric Tissue Spatial Transcriptomics from Serial Histology

### v1

@@ 1 Introduction
Spatial transcriptomics [Williams2022] links gene expression to tissue morphology and has transformed the way we study cellular states and disease-associated niches [McGregor2025, Kanemaru2023] in situ [Zhang2022]. Existing platforms (e.g., Visium, Xenium, CosMx) operate on 2D sections with a limited field of view, capturing only a thin slice of a much larger 3D tissue volume. This restricts our ability to probe vertical cell–cell communication, 3D tissue niches, and morphomolecular architectures that span multiple planes. Native 3D or thick-tissue in situ assays and 3D molecular atlases partially address this [Fang_2024, Sui2024-ct, doi:10.1126/science.adq2084], but remain constrained by capture depth and area, long protocols, specialized instrumentation, and high cost. Another approach is to densely section the tissue and perform ST on many serial sections [Oliveira2025, Mo2024], but this remains prohibitively expensive and labor-intensive to reconstruct full 3D volumes in routine studies.
@@
Within this emerging paradigm of H&E-based ST prediction, two main streams of work have appeared. The first focuses on whole-slide generative modeling with flow matching, exemplified by STFlow [HuangSTFlow], which models the joint distribution of all spots on a slide using local spatial attention and efficient slide-level encoders. The second centers on probabilistic assignment and deconvolution frameworks such as ST-Assign [stassign2023], which integrate scRNA-seq references with ST to infer cell-type mixtures and domains. While effective, these approaches leave open two key gaps for 3D reconstruction from serial histology: (i) most H&E to ST predictors are trained per section and do not explicitly enforce 3D coherence across serial stacks, leading to drift along the $z$-axis; and (ii) the start distribution used by flow-matching models is rarely tailored to the zero-inflated, overdispersed nature of gene counts  [Lopez2022, lopez2018scvi], placing undue burden on the denoiser early in the generative trajectory.
To address these gaps, we introduce HoloTea, a 3D H&E to ST leveraging flow matching [lipman2022flowmatching, liu2023rectifiedflow]. For each query spot on section $z$, we retrieve morphologically corresponding candidates on sections $z{\pm}1$ using $(x,y)$ proximity and feature-space cosine similarity, then inject this adjacent context into selected denoiser blocks via a lightweight ControlNet [zhang2023controlnet]. This design promotes anatomical continuity across sections without heavy preprocessing. Because flow matching decouples the probability path from the start distribution, we can instantiate biology-compatible priors: a learned ZINB prior using the scVI parameterization [lopez2018scvi, gayoso2022scvitools] and a 3D-consistent spatial–empirical prior constructed from neighboring sections. Compared to STFlow [HuangSTFlow], we extend flow matching based ST generation from single 2D sections to the serial-section regime, coupling FM stability and slide-level efficiency with explicit 3D conditioning and biology-aligned priors for transcript count data. In contrast to ST-Assign [stassign2023], which relies on external scRNA-seq references and global mixture inference, our H&E-only generator is fully reference-free and incurs a training and inference cost that is linear in the number of spots (local kNN graphs plus $O(Nm)$ inducing-point attention), naturally scaling to whole slides and serial stacks. An overview of the retrieval-conditioned flow-matching pipeline including $z{\pm}1$ neighbor search, adjacent-token aggregation, lightweight ControlNet conditioning, and ZINB-based priors is shown in Fig. 2.
@@ 2.1 H&E to Expression prediction
Large-scale H&E to ST models have advanced along three axes: stronger multi-resolution encoders, better supervision/benchmarks, and generative objectives. TRIPLEX aggregates cellular-, meso-, and slide-scale cues with an explicit fusion module and shows consistent gains on Visium datasets [Chung2024TRIPLEX]. SpaRED curates a 26-source benchmark with harmonized preprocessing and demonstrates that completing corrupted ST matrices significantly boosts downstream H&E to ST predictors [Mejia2024SpaRED]. Beyond purely discriminative regressors, recent generative approaches treat ST as a conditional generation problem: Stem leverages diffusion modeling to capture stochasticity in spot-level gene profiles given histology and reports SOTA across several datasets [Zhu2025DiffSGE]. Complementary alignment-based methods (e.g., DANet) emphasize dynamic cross-modal alignment to couple image features with gene targets while accounting for gene–gene dependencies [Wu2025DANet]. Meanwhile, multimodal contrastive formulations (e.g., mclSTExp) integrate spot context with transformer attention and image–omics contrastive learning [Min2024mclSTExp]. Finally, pathology foundation encoders (UNI/UNI2) pretrained on 100100M–200200M patches are increasingly used as robust backbones for H&E feature extraction in ST pipelines [UNI_NatMed2024, UNI_GitHub, UNI_HF].
@@
On the data and systems side, Open-ST introduces practical end-to-end protocols and tooling for subcellular ST in 2D/3D, including serial-section registration and 3D “virtual tissue block” reconstruction that lower the barrier to volumetric studies [OpenST_Cell_2024, OpenST_StarProt_2025, OpenST_Docs_2024]. For multi-slice analytics, SpaDo contributes cross-slice spatial domain detection, reference-based annotation, and large-scale clustering across platforms, providing infrastructure to reason about consistency and heterogeneity in stacks [Duan2024SpaDo]. In parallel, the field is moving toward larger multimodal corpora that pair WSIs with ST at scale: HEST-1k consolidates 1,2291{,}229 paired ST and H&E samples (2.1M expression–morphology pairs) and a library that standardizes ingestion and benchmarking, facilitating 3D/stack-aware evaluation and the use of strong pathology backbones for volumetric imputation [HEST2024, HEST_GitHub, HEST_NeurIPSProc].
@@ 3.2 Background
where $\mathbf{c}_{i}$ concatenates the image features $\mathbf{v}_{i}$, the coordinates $\mathbf{s}_{i}$, and a neighborhood summary $\mathrm{kNN}(i)$ (the indices or pooled features of the $k$ nearest spots to $i$ within its section). We minimize the supervised endpoint-regression loss $\mathcal{L}_{\mathrm{fm}}=\mathbb{E}_{i,t}\big[\|\hat{\mathbf{y}}_{i}-\mathbf{y}_{i}\|_{2}^{2}\big]$, following standard FM practice for displacement or straight paths [lipman2022flowmatching, liu2022rectifiedflow, esser2024rectified].
@@
A practical strength of FM is that the objective above does not depend on the form of $p_{0}$, so we can choose a starting prior that matches the domain to shorten the transport. STFlow adopts a ZINB prior for gene counts due to over-dispersion and excess zeros in ST data [HuangSTFlow]. Following the scVI  [lopez2018scvi, gayoso2022scvitools], for each gene $g$ and spot $i$ we use mean $\mu_{i,g}\!>\!0$, inverse-dispersion $\theta_{g}\!>\!0$, and zero-inflation probability $\pi_{i,g}\!\in\!(0,1)$, with Negative Binomial pmf
@@
Here $y$ is the gene-count random variable, $\Gamma(\cdot)$ is the gamma function, and the ZINB parameters may include library-size factors in $\mu_{i,g}$ and logit parameterization for $\pi_{i,g}$ as in scvi. We sample $\mathbf{Y}_{i}\sim\prod_{g=1}^{G}\mathrm{ZINB}(\mu_{i,g},\theta_{g},\pi_{i,g})$ and map to the model domain via $\mathbf{x}_{0,i}=\log(1+\mathbf{Y}_{i})$ . ZINB is standard for scRNA-seq and ST likelihoods [risso2018zinbwave] and is implemented in scvi-tools [gayoso2022scvitools].
@@
where $\mathbf{q}_{i},\mathbf{k}_{j},\mathbf{v}_{j}^{\text{(val)}}$ are the query, key, and value vectors for spots $i$ and $j$, $d$ is the per-head key/query dimensionality, $\mathcal{N}(i)$ is the k-NN set of $i$, $\alpha_{ij}$ are attention weights, and $\mathbf{h}_{i}$ is the aggregated representation. This graph-attention view follows established formulations [velickovic2018gat] and can be equipped with inducing-point self-attention for low-rank global mixing [lee2019settransformer]; see also the whole-slide, locally attentive denoiser in STFlow [HuangSTFlow].
@@ 4 Method
Our method contains three main components: (i) conditioning from adjacent spots on the closest slides together with a prior over gene–gene relations in training dataset; (ii) applying conditioning and $z$-variation through ControlNet [zhang2023controlnet]; and (iii) building hierarchical feature selection efficiently via Global set attention (GSA) [lee2019settransformer].
@@ 4.1 Adjacent-section conditioning
where $\mathrm{Proj}_{\ell}(\cdot)$ linearly matches $\mathbf{u}_{i}$ to the block’s hidden size, and $\alpha_{\ell}(t)$ is a warm-up schedule in $t$ to avoid early over-steering; a sharp variant is used to preserve high-frequency detail. This follows the conditioning principle of ControlNet [zhang2023controlnet] while adapting it to spot-wise conditioning.
@@ 4.2 $z$-aware priors
where $\bm{\mu}_{i}$ are per-gene means, $\bm{\theta}_{i}$ inverse-dispersions, and $\bm{\pi}_{i}$ zero-inflation probabilities (scVI parameterization [lopez2018scvi, gayoso2022scvitools]; see ZINB in Sec. 3.2). The pretraining minimizes the negative log-likelihood of the ZINB model on the training expressions:
@@ 4.3 Global set attention (GSA)
To inject global context without quadratic cost in the number of spots $N$, we use a two-stage attention block inspired by the Induced Set Attention Block (ISAB) [lee2019settransformer]. Let $\mathbf{X}\!\in\!\mathbb{R}^{N\times d}$ be spot tokens (per spot: current state $\mathbf{x}_{t,i}$, H&E features, time/positional embeddings) and let $\mathbf{S}\!\in\!\mathbb{R}^{m\times d}$ be a small set of inducing tokens ($m\!\ll\!N$), learned and shared across batches. We use multi-head attention $\mathrm{MHA}(\mathbf{Q},\mathbf{K},\mathbf{V}){=}\mathrm{Concat}_{h}\!\big(\mathrm{softmax}(\mathbf{Q}\mathbf{K}^{\top}/\sqrt{d_{h}})\mathbf{V}\big)\mathbf{W}^{O}$ with $h$ heads and per-head width $d_{h}{=}d/h$.
@@ 5.1 Experiments settings
We evaluate on two public benchmarks: the HER2 multi-section breast cancer dataset [Andersson2021HER2NatComm] and the ST-Data breast cancer cohort introduced with ST-Net [He2020STNet]. We follow the authors’ preprocessing and adopt section-level holdouts for testing (entire heldout sections), mirroring typical cross-section protocols described for these datasets. Moreover, we benchmark on the in-house Embryo dataset (to be released soon in a separate manuscript), it contains a 3D stack of serial embryo sections profiled with 10X Visium HD, yielding dense spot-level transcript counts co-registered with H&E whole-slide images (and optional auxiliary imaging channels). The stack provides a multi-modal volume (H&E ++ ST) across 52 slides of z-stack of H&E.
@@
All experiments use PyTorch with mixed precision, Adam optimizer [kingma2015adam], and a single NVIDIA H100 (80 GB). H&E tiles are encoded by the UNI pathology foundation model [Chen2024UNI_NatMed]. Our ZINB decoder follows the scVI parameterization and NLL [lopez2018scvi, gayoso2022scvitools], taking as input either the STFlow latent representation or features from the frozen UNI pathology encoder [Chen2024UNI_NatMed]. In the UNI baseline, we keep the UNI backbone fixed and train only this ZINB head, yielding an image-only encoder–decoder model without any flow-based components. Flow training uses displacement/rectified flow matching [lipman2022flowmatching, liu2022rectifiedflow], and our ControlNet-style conditioning for $z$-aware guidance follows [zhang2023controlnet]. The training follows Algorithm 1.
For ASIGN 3D [Zhu_2025_CVPR_ASIGN], we re-ran the method on the HER2 and ST imputation tasks using the normalization described in [HuangSTFlow], enabling a fair comparison of MAE and MSE. On the in-house embryo dataset, the original ASIGN 3D configuration resulted in out-of-memory (OOM) errors even with a batch size of one; therefore, we reduced the model size by removing one graph neural network layer, and all reported results for this dataset correspond to that reduced configuration. Hyperparameters are listed in Appendix LABEL:app:experiments_details.
@@
We report results on two panels: (i) Top-250 HVG selected with Scanpy’s pp.highly_variable_genes (Seurat-flavor) [wolf2018scanpy, Stuart2019SeuratV3]; (ii) Custom gene set, built by detecting spatially variable genes (SVGs) per section via SPARK-X [sun2021sparkx], intersecting the top-$K$ SVGs within sliding windows of five consecutive sections, and taking the union across windows to stabilize across-$z$ morphology; (iii) marker genes for 106 anatomical regions were identified by performing differential gene expression analysis. Selecting the top two marker genes for each region yielded a set of 211 distinct genes.
@@ 5.2 Gene Expression Imputation
We perform 3D spatial imputation as generating a whole slide gene expression from serial H&E sections, given ground-truth ST on only a subset of sections. We adopt two complementary splits. (i) Following the ASIGN 3D-ST imputation task [Zhu_2025_CVPR_ASIGN], each 3D sample provides a single ST-labeled section used for supervision, while all remaining sections in the same stack are ST-unlabeled and serve as imputation targets. Then models are trained across samples and evaluated on entire held-out sections without using their ST during training. (ii) Even-slice hold-out. Within each 3D stack, we hold out every even-index section ($z{=}2,4,\dots$) as test targets and train on the odd-index sections ($z{=}1,3,\dots$), using only their ST labels. This split probes across-$z$ generalization under a fixed anatomy and avoids reliance on any single ‘known’ slice.
We conduct both splits on two public multi-section breast datasets, HER2 [Andersson2021HER2NatComm] and ST-Data [He2020STNet]. For our in-house embryo dataset, we use the even-slice hold-out split to simulate missing-through-plane acquisition at scale. We report the results for the whole-slide imputation task in Tab. 1 and Tab. 2. HoloTea outperforms other methods across all three datasets and splits. ASIGN performs poorly on Embryo since the memory scales with number of slides and number of spots per slide, we discuss the details in Sec. 5.4.

### v2

@@ 1 Introduction
Spatial transcriptomics [40] links gene expression to tissue morphology and has transformed the way we study cellular states and disease-associated niches [29, 15] in situ [44]. Existing platforms (e.g., Visium, Xenium, CosMx) operate on 2D sections with a limited field of view, capturing only a thin slice of a much larger 3D tissue volume. This restricts our ability to probe vertical cell–cell communication, 3D tissue niches, and morphomolecular architectures that span multiple planes. Native 3D or thick-tissue in situ assays and 3D molecular atlases partially address this [7, 37, 8], but remain constrained by capture depth and area, long protocols, specialized instrumentation, and high cost. Another approach is to densely section the tissue and perform ST on many serial sections [33, 32], but this remains prohibitively expensive and labor-intensive to reconstruct full 3D volumes in routine studies.
@@
Within this emerging paradigm of H&E-based ST prediction, two main streams of work have appeared. The first focuses on whole-slide generative modeling with flow matching, exemplified by STFlow [12], which models the joint distribution of all spots on a slide using local spatial attention and efficient slide-level encoders. The second centers on probabilistic assignment and deconvolution frameworks such as ST-Assign [10], which integrate scRNA-seq references with ST to infer cell-type mixtures and domains. While effective, these approaches leave open two key gaps for 3D reconstruction from serial histology: (i) most H&E to ST predictors are trained per section and do not explicitly enforce 3D coherence across serial stacks, leading to drift along the $z$-axis; and (ii) the start distribution used by flow-matching models is rarely tailored to the zero-inflated, overdispersed nature of gene counts  [26, 25], placing undue burden on the denoiser early in the generative trajectory.
To address these gaps, we introduce HoloTea, a 3D H&E to ST leveraging flow matching [22, 24]. For each query spot on section $z$, we retrieve morphologically corresponding candidates on sections $z{\pm}1$ using $(x,y)$ proximity and feature-space cosine similarity, then inject this adjacent context into selected denoiser blocks via a lightweight ControlNet [45]. This design promotes anatomical continuity across sections without heavy preprocessing. Because flow matching decouples the probability path from the start distribution, we can instantiate biology-compatible priors: a learned ZINB prior using the scVI parameterization [25, 9] and a 3D-consistent spatial–empirical prior constructed from neighboring sections. Compared to STFlow [12], we extend flow matching based ST generation from single 2D sections to the serial-section regime, coupling FM stability and slide-level efficiency with explicit 3D conditioning and biology-aligned priors for transcript count data. In contrast to ST-Assign [10], which relies on external scRNA-seq references and global mixture inference, our H&E-only generator is fully reference-free and incurs a training and inference cost that is linear in the number of spots (local kNN graphs plus $O(Nm)$ inducing-point attention), naturally scaling to whole slides and serial stacks. An overview of the retrieval-conditioned flow-matching pipeline including $z{\pm}1$ neighbor search, adjacent-token aggregation, lightweight ControlNet conditioning, and ZINB-based priors is shown in Fig. 2.
@@ 2.1 H&E to Expression prediction
Large-scale H&E to ST models have advanced along three axes: stronger multi-resolution encoders, better supervision/benchmarks, and generative objectives. TRIPLEX aggregates cellular-, meso-, and slide-scale cues with an explicit fusion module and shows consistent gains on Visium datasets [3]. SpaRED curates a 26-source benchmark with harmonized preprocessing and demonstrates that completing corrupted ST matrices significantly boosts downstream H&E to ST predictors [30]. Beyond purely discriminative regressors, recent generative approaches treat ST as a conditional generation problem: Stem leverages diffusion modeling to capture stochasticity in spot-level gene profiles given histology and reports SOTA across several datasets [48]. Complementary alignment-based methods (e.g., DANet) emphasize dynamic cross-modal alignment to couple image features with gene targets while accounting for gene–gene dependencies [42]. Meanwhile, multimodal contrastive formulations (e.g., mclSTExp) integrate spot context with transformer attention and image–omics contrastive learning [31]. Finally, pathology foundation encoders (UNI/UNI2) pretrained on 100100M–200200M patches are increasingly used as robust backbones for H&E feature extraction in ST pipelines [16, 20, 28].
@@
On the data and systems side, Open-ST introduces practical end-to-end protocols and tooling for subcellular ST in 2D/3D, including serial-section registration and 3D “virtual tissue block” reconstruction that lower the barrier to volumetric studies [46, 27, 34]. For multi-slice analytics, SpaDo contributes cross-slice spatial domain detection, reference-based annotation, and large-scale clustering across platforms, providing infrastructure to reason about consistency and heterogeneity in stacks [5]. In parallel, the field is moving toward larger multimodal corpora that pair WSIs with ST at scale: HEST-1k consolidates 1,2291{,}229 paired ST and H&E samples (2.1M expression–morphology pairs) and a library that standardizes ingestion and benchmarking, facilitating 3D/stack-aware evaluation and the use of strong pathology backbones for volumetric imputation [14, 19, 13].
@@ 3.2 Background
where $\mathbf{c}_{i}$ concatenates the image features $\mathbf{v}_{i}$, the coordinates $\mathbf{s}_{i}$, and a neighborhood summary $\mathrm{kNN}(i)$ (the indices or pooled features of the $k$ nearest spots to $i$ within its section). We minimize the supervised endpoint-regression loss $\mathcal{L}_{\mathrm{fm}}=\mathbb{E}_{i,t}\big[\|\hat{\mathbf{y}}_{i}-\mathbf{y}_{i}\|_{2}^{2}\big]$, following standard FM practice for displacement or straight paths [22, 23, 6].
@@
A practical strength of FM is that the objective above does not depend on the form of $p_{0}$, so we can choose a starting prior that matches the domain to shorten the transport. STFlow adopts a ZINB prior for gene counts due to over-dispersion and excess zeros in ST data [12]. Following the scVI  [25, 9], for each gene $g$ and spot $i$ we use mean $\mu_{i,g}\!>\!0$, inverse-dispersion $\theta_{g}\!>\!0$, and zero-inflation probability $\pi_{i,g}\!\in\!(0,1)$, with Negative Binomial pmf
@@
Here $y$ is the gene-count random variable, $\Gamma(\cdot)$ is the gamma function, and the ZINB parameters may include library-size factors in $\mu_{i,g}$ and logit parameterization for $\pi_{i,g}$ as in scvi. We sample $\mathbf{Y}_{i}\sim\prod_{g=1}^{G}\mathrm{ZINB}(\mu_{i,g},\theta_{g},\pi_{i,g})$ and map to the model domain via $\mathbf{x}_{0,i}=\log(1+\mathbf{Y}_{i})$ . ZINB is standard for scRNA-seq and ST likelihoods [35] and is implemented in scvi-tools [9].
@@
where $\mathbf{q}_{i},\mathbf{k}_{j},\mathbf{v}_{j}^{\text{(val)}}$ are the query, key, and value vectors for spots $i$ and $j$, $d$ is the per-head key/query dimensionality, $\mathcal{N}(i)$ is the k-NN set of $i$, $\alpha_{ij}$ are attention weights, and $\mathbf{h}_{i}$ is the aggregated representation. This graph-attention view follows established formulations [39] and can be equipped with inducing-point self-attention for low-rank global mixing [21]; see also the whole-slide, locally attentive denoiser in STFlow [12].
@@ 4 Method
Our method contains three main components: (i) conditioning from adjacent spots on the closest slides together with a prior over gene–gene relations in training dataset; (ii) applying conditioning and $z$-variation through ControlNet [45]; and (iii) building hierarchical feature selection efficiently via Global set attention (GSA) [21].
@@ 4.1 Adjacent-section conditioning
where $\mathrm{Proj}_{\ell}(\cdot)$ linearly matches $\mathbf{u}_{i}$ to the block’s hidden size, and $\alpha_{\ell}(t)$ is a warm-up schedule in $t$ to avoid early over-steering; a sharp variant is used to preserve high-frequency detail. This follows the conditioning principle of ControlNet [45] while adapting it to spot-wise conditioning.
@@ 4.2 $z$-aware priors
where $\bm{\mu}_{i}$ are per-gene means, $\bm{\theta}_{i}$ inverse-dispersions, and $\bm{\pi}_{i}$ zero-inflation probabilities (scVI parameterization [25, 9]; see ZINB in Sec. 3.2). The pretraining minimizes the negative log-likelihood of the ZINB model on the training expressions:
@@ 4.3 Global set attention (GSA)
To inject global context without quadratic cost in the number of spots $N$, we use a two-stage attention block inspired by the Induced Set Attention Block (ISAB) [21]. Let $\mathbf{X}\!\in\!\mathbb{R}^{N\times d}$ be spot tokens (per spot: current state $\mathbf{x}_{t,i}$, H&E features, time/positional embeddings) and let $\mathbf{S}\!\in\!\mathbb{R}^{m\times d}$ be a small set of inducing tokens ($m\!\ll\!N$), learned and shared across batches. We use multi-head attention $\mathrm{MHA}(\mathbf{Q},\mathbf{K},\mathbf{V}){=}\mathrm{Concat}_{h}\!\big(\mathrm{softmax}(\mathbf{Q}\mathbf{K}^{\top}/\sqrt{d_{h}})\mathbf{V}\big)\mathbf{W}^{O}$ with $h$ heads and per-head width $d_{h}{=}d/h$.
@@ 5.1 Experiments settings
We evaluate on two public benchmarks: the HER2 multi-section breast cancer dataset [1] and the ST-Data breast cancer cohort introduced with ST-Net [11]. We follow the authors’ preprocessing and adopt section-level holdouts for testing (entire heldout sections), mirroring typical cross-section protocols described for these datasets. Moreover, we benchmark on the in-house Embryo dataset (to be released soon in a separate manuscript), it contains a 3D stack of serial embryo sections profiled with 10X Visium HD, yielding dense spot-level transcript counts co-registered with H&E whole-slide images (and optional auxiliary imaging channels). The stack provides a multi-modal volume (H&E ++ ST) across 52 slides of z-stack of H&E.
@@
All experiments use PyTorch with mixed precision, Adam optimizer [17], and a single NVIDIA H100 (80 GB). H&E tiles are encoded by the UNI pathology foundation model [2]. Our ZINB decoder follows the scVI parameterization and NLL [25, 9], taking as input either the STFlow latent representation or features from the frozen UNI pathology encoder [2]. In the UNI baseline, we keep the UNI backbone fixed and train only this ZINB head, yielding an image-only encoder–decoder model without any flow-based components. Flow training uses displacement/rectified flow matching [22, 23], and our ControlNet-style conditioning for $z$-aware guidance follows [45]. The training follows Algorithm 1.
For ASIGN 3D [47], we re-ran the method on the HER2 and ST imputation tasks using the normalization described in [12], enabling a fair comparison of MAE and MSE. On the in-house embryo dataset, the original ASIGN 3D configuration resulted in out-of-memory (OOM) errors even with a batch size of one; therefore, we reduced the model size by removing one graph neural network layer, and all reported results for this dataset correspond to that reduced configuration. Hyperparameters are listed in Appendix S2.
@@
We report results on two panels: (i) Top-250 HVG selected with Scanpy’s pp.highly_variable_genes (Seurat-flavor) [41, 36]; (ii) Custom gene set, built by detecting spatially variable genes (SVGs) per section via SPARK-X [38], intersecting the top-$K$ SVGs within sliding windows of five consecutive sections, and taking the union across windows to stabilize across-$z$ morphology; (iii) marker genes for 106 anatomical regions were identified by performing differential gene expression analysis. Selecting the top two marker genes for each region yielded a set of 211 distinct genes.
@@ 5.2 Gene Expression Imputation
We perform 3D spatial imputation as generating a whole slide gene expression from serial H&E sections, given ground-truth ST on only a subset of sections. We adopt two complementary splits. (i) Following the ASIGN 3D-ST imputation task [47], each 3D sample provides a single ST-labeled section used for supervision, while all remaining sections in the same stack are ST-unlabeled and serve as imputation targets. Then models are trained across samples and evaluated on entire held-out sections without using their ST during training. (ii) Even-slice hold-out. Within each 3D stack, we hold out every even-index section ($z{=}2,4,\dots$) as test targets and train on the odd-index sections ($z{=}1,3,\dots$), using only their ST labels. This split probes across-$z$ generalization under a fixed anatomy and avoids reliance on any single ‘known’ slice.
We conduct both splits on two public multi-section breast datasets, HER2 [1] and ST-Data [11]. For our in-house embryo dataset, we use the even-slice hold-out split to simulate missing-through-plane acquisition at scale. We report the results for the whole-slide imputation task in Tab. S1 and Tab. 2. HoloTea outperforms other methods across all three datasets and splits. ASIGN performs poorly on Embryo since the memory scales with number of slides and number of spots per slide, we discuss the details in Sec. 5.4.